KyledaGreat:

A certain species of deer is to be introduced into a forest, and wildlife experts estimate the population will grow to P(t)=(727)3^t/4, where t represents the number of years from the time of introduction. Step 1 of 2 : What is the tripling-time for this population of deer?

KyledaGreat:

@vocaloid

KyledaGreat:

sorry my computr acted up when i tagged you once

Vocaloid:

P(t)=(727)3^(t/4) let's just look at this part: 3^(t/4) notice how, when t = 4, the expression becomes 3^(4/4) = 1 so when t = 4, P(t)=(727)3^(t/4) = (727) * 3 * 1 so we conclude that the tripling time is t = 4 yrs

KyledaGreat:

How long will it take for the population to reach 6543 deer, according to this model?

Vocaloid:

P(t) = 6543 = (727)3^(t/4) solve for t

KyledaGreat:

8

Vocaloid:

good, so 8 yrs

KyledaGreat:

okay